Trace the curve y =
\sqrt[3]{ {x}^{2} - 1}
3

x
2
−1


and state all the properties you use to trace it
1
Expert's answer
2020-03-09T13:57:34-0400

Solution:

"~\\y=\sqrt[3]{x^2-1}"


1.Thej domain :"~x\in(-\infty,+\infty)"


2. Intersection points with coordinate axes:

"Ox: y=0\to x^2-1=0 \to x_1=1, x_2=-1\to A(1,0), B(-1,0)\\
 Oy:x=0\to y^3=-1\to y=-1\to C(0,-1)"

the graph is symmetric about the axis Oy


4. The function is continuous


5. We find the critical points: "y'=\frac{1}{3}(x^2-1)^{-\frac{2}{3}}\cdot2x\to x=0, x=1,x=-1"


"x\in(-\infty,-1)\cup(-1,0), y' \lt 0\to y\searrow\\ x\in(0,1)\cup(1,\infty), y' \gt 0 \to y\nearrow\\"


Then "x=0"  is the minimum point "y=-1,x=0"


6. Find the inflection points:

"y''=\frac{2}{3}(x^2-1)^{-\frac{2}{3}}+\frac{2}{3}x\cdot(-\frac{2}{3})(x^2-1)^{-\frac{5}{3}} ="

"\frac{2}{3}(x^2-1)^{-\frac{5}{3}}(x^2-1-\frac{4}{3}x^2) =\frac{2}{3}(x^2-1)^{-\frac{5}{3}}(-\frac{1}{3}x^2-1)\\~\\ x\in(-\infty,-1)\cup(1,+\infty)\to y'' \lt 0\\ x\in (-1,1)\to y'' \gt 0"



7. There are no vertical asymptotes horizontal asymptote:"y=kx+b~~~\to~~~ k=\lim_{x\to\infty} \frac{\sqrt[3]{x^2-1}}{x}=0\\ b=\lim_{x\to\infty} (\sqrt[3]{x^2-1}-0\cdot x)=\infty"

there are no horizontaltical asymptotes


8. Intervals for the range:

"y \gt 0 \to x^2-1 \gt 0,~ x^2 \gt 1 \to x\in(-\infty;-1)\cap (1;\infty)\\
 y \lt 0\to x\in(-1;1)"


9. Behavior at infinity: "\begin{bmatrix}
 x\to-\infty, y\to+\infty\\
x\to+\infty, y\to+\infty
\end{bmatrix}"